IF LEAH ID 6 YEARS OLD THAN HER SISTER, SUE, ADN JOHN IS 5 YEARS OLDER THAN LEAH, AND THE TOTAL OF THEIR AGES IS 42. THEN HOW OLD IS SUE?
Answer by edjones(7311)

Let L=leah's age. L-6=Sue's age L+5=John's age
.
L+L-6+L+5=42
3L=43
L=14 1/3
8 1/3=Sue's age
19 1/3=John's age
---------
41 3/3 =42
.
Ed